OpenStudy (anonymous):

What are the restrictions on squaring both sides of an equation. Please provide a detailed discussion.

OpenStudy (anonymous):

I can't think of any restrictions on squaring both sides of an equation.

OpenStudy (anonymous):

Can you square the equation |3|=x I think if you do that you will loose solutions.

OpenStudy (anonymous):

yes, you are right.

OpenStudy (anonymous):

But I know that that knowledge is not enough. Can you please provide me some more of such things

OpenStudy (anonymous):

I mean a little details, and when not to square etc. etc

OpenStudy (anonymous):

I would, but I am very sleepy now. I'll think of some and come back. This is interesting stuff.

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

you should ask INewton and Lokisan. I think they will be able to help you the best.

OpenStudy (anonymous):

Ok I will do that

OpenStudy (anonymous):

Glad that you are here

OpenStudy (anonymous):

Could you help please

OpenStudy (anonymous):

There are no restrictions.

OpenStudy (anonymous):

What is causing you to think that?

OpenStudy (anonymous):

You see, I often get more solutions than what I am expected to get, when I square both sides

OpenStudy (anonymous):

For example If I start with sqrt 2 + sqrt 5 =x sqrt 2 =x - sqrt 5

OpenStudy (anonymous):

and then square both sides and go on, I will end up with sqrt 2 +- sqrt 5

OpenStudy (anonymous):

I'm not following. If you square both sides, you should get\[(\sqrt{2})^2=(x-\sqrt{5})^2 \rightarrow 2=x^2-2\sqrt{5}x+5\]

OpenStudy (anonymous):

So now if we use the formula for quadratic equation

OpenStudy (anonymous):

We get two things

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You're getting two answers now because you've asked a different question.

OpenStudy (anonymous):

"different question"?

OpenStudy (anonymous):

Solving your first equation is not the same as solving the squared one.

OpenStudy (anonymous):

yes, i think lokisan is right. you asked a different question when you take square root on both sides

OpenStudy (anonymous):

I see what you're thinking, but you needn't worry :)

OpenStudy (anonymous):

You have two choices going on here: 1) take your expression as is and solve 2) operate on your expression with an operator that maps this thing to something else and solve for that When you square, you're doing the latter,

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So does that mean I shouldn't square any equation

OpenStudy (anonymous):

I know you're not convinced, but just think about it for a while.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Why did you think you had to square it anyway?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Just working with irrational numbers

OpenStudy (anonymous):

I was trying to prove that the sum of these two irrational numbers will lead to one irrational number

OpenStudy (anonymous):

You can't get rid of irrational sums by squaring!

OpenStudy (anonymous):

Oh...see above.

OpenStudy (anonymous):

I have another problem here sqrt (x+1)-sqrt (x-1)=sqrt (4x-1) Now if I use wolframalfa, I find no solution But if I square both sides, I do find

OpenStudy (anonymous):

This made me question, "what are the restrictions..........."

OpenStudy (anonymous):

Well, I haven't checked it, but if it has no solutions, it just means there are no x-values that will satisfy that expression; it's stuffing up because there are no possible x-values that everything can share that won't make the radicand (thing under sqrt) negative.

OpenStudy (anonymous):

Even if you square, you're going to end up with a situation where no x-values are available to make that statement true. http://www.wolframalpha.com/input/?i=%5Bsqrt+%28x%2B1%29-sqrt+%28x-1%29%5D^2%3Dsqrt+%284x-1%29^2

OpenStudy (anonymous):

But if you try squaring them, you will find that we get a solution

OpenStudy (anonymous):

And you will have to square twice

OpenStudy (anonymous):

nope, wolfram alfa says no solutions exist.

OpenStudy (anonymous):

If you have time, then I can square both sides, and solve it here

OpenStudy (anonymous):

I will do it manually, on paper.

OpenStudy (anonymous):

Yes, that is what I am talking of

OpenStudy (anonymous):

yes, so solve it here, iamignorant/

OpenStudy (anonymous):

sqrt(x+1)-sqrt(x-1)=sqrt(4x-1)

OpenStudy (anonymous):

x+1+x-1-2sqrt(x+1)(x-1)=4x-1

OpenStudy (anonymous):

2x-2sqrt(x+1)(x-1)=4x-1

OpenStudy (anonymous):

-2sqrt(x+1)(x-1)=2x-1

OpenStudy (anonymous):

You do get a solution, x=5/4, but when you test it back in the original equation, the left-hand side is 1, while the right-hand side is 2. This is a contradiction.

OpenStudy (anonymous):

Yes exactly

OpenStudy (anonymous):

But how do I know

OpenStudy (anonymous):

I mean will I have to plug in every value after solving every equation of the same type

OpenStudy (anonymous):

You ended up at a contradiction assuming this thing had a solution. The contradiction shows the assumption to be false (reductio ad absurdum).

OpenStudy (anonymous):

In general: You can square an equation and solve as normal, but you must ALWAYS check the solutions in the original, as squaring can produce spurious roots.

OpenStudy (anonymous):

Just what I said.

OpenStudy (anonymous):

Does that mean I must always check back?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now what to say......

OpenStudy (anonymous):

How about, "I get it"?

OpenStudy (anonymous):

sorry?

OpenStudy (anonymous):

No...you just learnt something.

OpenStudy (anonymous):

I am sorry, but I am still not getting you

OpenStudy (anonymous):

Sorry I didn't read the whole thing (after you said "There are no restrictions"..., Lokisan). Regardless, I would not phrase it as you have (with reductio ad absurdum etc...)

OpenStudy (anonymous):

Well, it's late, so I'm not really bothered with semantics.

OpenStudy (anonymous):

Are you satisfied, Iam?

OpenStudy (anonymous):

I will have to be, no other option

OpenStudy (anonymous):

I have to go - just think about it...

OpenStudy (anonymous):

:Sure I will

OpenStudy (anonymous):

Geeeeeez that guy's a buzzkill.

OpenStudy (anonymous):

Whatever, he has helped me a lot since the time he joined openstudy

OpenStudy (anonymous):

And also is a smart man

OpenStudy (anonymous):

That doesn't mean I am trying to offend you Mr Newton

OpenStudy (anonymous):

My comment was somewhat tongue-in-cheek. But, how smart?

OpenStudy (anonymous):

Actually I am ignorant, and any comment from my side will simply be a ridiculous one

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Any how

OpenStudy (anonymous):

Mr Newton, this thing of plugging final solutions, and checking them

OpenStudy (anonymous):

What about it?

OpenStudy (anonymous):

Should it be done always, or only when we are dealing with square roots

OpenStudy (anonymous):

There is no need 'normally' if your steps all follow. Only when you have needed to square the equation (i.e. basically when you have roots etc)

OpenStudy (anonymous):

You mean, when I am searching for roots, only then I need to actually do these things

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

I meant square roots. Roots of the equation would not (normally) have to be checked.

OpenStudy (anonymous):

If you notice some of our conversation at the beginning, you will notice, that once Mr Lucas said that I have actually changed my question by squaring both sides

OpenStudy (anonymous):

This was the question

OpenStudy (anonymous):

If I start with sqrt 2 + sqrt 5 =x sqrt 2 =x - sqrt 5

OpenStudy (anonymous):

I squared both sides and then I solved it

OpenStudy (anonymous):

I ended up with sqrt 2 +- sqrt 5

OpenStudy (anonymous):

So Mr Lucas commented, that I actually changed the question by squaring it

OpenStudy (anonymous):

Could you please provide me a little more explanation

OpenStudy (anonymous):

I mean where I changed it, and how did you all guess, that I changed it

OpenStudy (anonymous):

OK. Consider the follow equations: \[\sqrt{3x^2+1} + \sqrt{x} - 2x -1 = 0 \] \[\sqrt{3x^2+1} -2\sqrt{x} +x - 1 = 0\] \[\sqrt{3x^2+1} -2\sqrt{x} -x + 1 = 0\] To solve these, you will need to square them (once if you split it up smartly onto each side it smart, but maybe twice if you don't). This will give the same 3 roots for all equations (which it is left for you to find if you are bored). However, they each have different solutions. Why? BECAUSE SQUARING CAN (but will not always) PRODUCE SPURIOUS ROOTS, which must be checked and rejected. The algebraic manipulation (in this case, squaring), has changed what you are trying to solve, and given extra roots.

OpenStudy (anonymous):

If you only read one part, read the last paragraph. It will not ALWAYS change the values that solve it, but it can so they MUST be checked.

OpenStudy (anonymous):

Have a nice day.